# Matchings and loose cycles in the semirandom hypergraph model

### v1

@@ 3. Outline, and lemmas on an auxiliary bipartite graph
Let $\rho=\frac{1-c}{c}$. The outneighbours of each node are chosen uniformly with replacement from the other part. So, the indegree of a node $v\in\mathcal{A}$ is distributed as $\operatorname{Bin}(k\left|{\mathcal{B}}\right|,1/\left|{\mathcal{A}}\right|)$, with expectation $k\left|{\mathcal{B}}\right|/\left|{\mathcal{A}}\right|\leq\rho k$. The same bound holds for $v\in\mathcal{B}$ since $\left|{\mathcal{A}}\right|\leq\left|{\mathcal{B}}\right|$. For a given set $S$ of size $\delta N$, $Z=\operatorname{vol}(S)$ is a sum of the two types of binomials, both independent, so $Z$ is dominated by a binomial random variable $Z^{\prime}$ with $\mathbb{E}Z^{\prime}=\rho k\delta N$.
Let $\rho=\frac{1-c}{c}$ and let $\eta=\eta_{c,k}(\delta)$. Apply Chernoff’s inequality $\Pr(Z\geq\mathbb{E}Z+t)\leq\exp\left(-\frac{t^{2}}{2(\mathbb{E}Z+t/3)}\right)$ [15, eq. (2.5)] to $\Pr(Z^{\prime}\geq\eta N)$. This means taking $t=\eta N-\mathbb{E}Z^{\prime}=(r-1)\rho k\delta N$, where $r=\eta/(\rho k\delta)\geq 2$ by definition of $\eta_{c,k}(\delta)$. Then,
$\Pr(Z^{\prime}\geq\eta N)$ $\leq\exp\left(-\frac{(r-1)^{2}(\mathbb{E}Z^{\prime})^{2}}{2(1+\frac{r-1}{3})\mathbb{E}Z^{\prime}}\right)\;\leq\;\exp\left(-\frac{(r/2)^{2}}{2(\frac{2r}{3})}\mathbb{E}Z^{\prime}\right)\;=\;\exp\Big{(}-\tfrac{3}{16}r\cdot\rho\delta kN\Big{)}.$
@@
$\binom{N}{\delta N}\Pr(Z^{\prime}\geq\eta N)$ $\leq\exp\Big{(}\delta N\cdot\Big{(}\ln(e/\delta)-\frac{3}{16}r\rho k\Big{)}\Big{)}$
@@
Given $k\geq 10$ and $\rho\leq 1/100$. Let $\mathcal{D}$ be a uniformly random $k$-out bipartite digraph on parts $\mathcal{A}$ and $\mathcal{B}$. Consider the subgraph $\mathcal{D}^{\prime}$ induced by a pair of subsets $\mathcal{A}^{\prime}\subseteq\mathcal{A}$, $\mathcal{B}^{\prime}\subseteq\mathcal{B}$, where
@@
Then, a.a.s. in $\left|{\mathcal{A}}\right|$, every such induced subgraph $\mathcal{D}^{\prime}$ contains a perfect matching.
Note that the asymptotic statement in terms of $\left|{\mathcal{A}}\right|$ is equivalent to one in terms of $\left|{\mathcal{B}}\right|$, $\left|{\mathcal{A}^{\prime}}\right|$, or $\left|{\mathcal{B}^{\prime}}\right|$. The hypothesis $k\geq 10$ can be weakened to $k\geq 7$, at the expense of a slightly more complicated proof, and likely further.
@@
the last inequality follows from the hypothesis on $\rho$ and is used in the next calculation. The same holds for $\mathcal{B}^{\prime}$. Also, $\left|{\mathcal{A}}\right|$ and $\left|{\mathcal{B}}\right|$ are both $\Theta(n)$, and henceforth we work asymptotically in $n$.
@@ 4.2.1. Phase 2a
(Read the label as M for matching, 2a for the phase, and C for “condition”, and in the next part, A for “action”.)
@@ 4.2.3. Analysis
The number of inneighbours of these nodes is bounded by Lemma 9. The lemma’s $N$ is $\left|{A_{1}}\right|+\left|{B_{1}}\right|>\frac{2}{3}n$ by (4.2) and (4.3), so $10\epsilon_{D}n$ represents at most a $15\epsilon_{D}$ fraction of all nodes; we take the lemma’s $\delta$ equal to $15\epsilon_{D}$. The two parts $A_{1}$ and $B_{1}$ are of nearly equal size, so we take the lemma’s $c$ as $\frac{1}{3}$. By the lemma, the number of inneighbours is at most $\eta_{1/3,k}(\epsilon_{D})N<\eta_{1/3,k}(\epsilon_{D})n$.
In the paragraph above, we reasoned that the parts are each of size about $\frac{1}{3}n$, and that (with generous allowance for the small difference) each is of size at least $c=1/3$ their combined size. There will be several such cases, so let us treat this one more carefully as an example. Each size is $\frac{1}{3}n+O(\epsilon_{D})n$, where the $O(\cdot)$ term indicates a value, positive or negative, bounded in absolute value by some constant such as 15 that we could work out by tracing calculations such as (4.1) and (4.2). We omit these calculations and similar ones, as we will take $\epsilon_{D}$ small enough that these $O(\cdot)$ terms are tiny, much less than the generous margins we will allow. More to the point, forgetting the constants calculated here, it is clear that $\epsilon_{D}$ can be chosen sufficiently small to make it so.
Each inneighbour has at most $k$ outneighbours, so the number of outneighbours is at most $k\>\eta_{1/3,k}(\epsilon_{D})n$. We would like this always to be small compared with $\left|{A}\right|$ and $\left|{B}\right|$, both of which will always be very nearly $\frac{1}{3}n$, and safely at least $\frac{1}{4}n$.
@@
This dictates our parameter choices. We will contrive that $k\>\eta_{1/3,k}(\epsilon_{D})\leq\frac{1}{101}\cdot\frac{1}{4}$, ensuring that the size of $Q$, including $Q_{1}$, satisfies $|{Q}|\leq\frac{1}{100}\min\left\{\left|{A_{0}}\right|,\left|{B_{0}}\right|\right\}$ at all times. We take $k=10$ for purposes of Lemma 10. We find that $\delta=$0.000\hskip 1.30005pt015$$ suffices, giving $\eta_{10,1/3}(\delta)<$0.001$<1/404$ as desired. (The value of $\eta$, defined in Lemma 9, is determined by the first term in the max; the second term, $40\>\delta$, is much smaller.) Since we set $\delta=15\epsilon_{D}$, taking $\epsilon_{D}=$0.000\hskip 1.30005pt001$$ suffices.
Looking ahead, we will invoke Lemma 10 to prove existence of a perfect matching in $D$. That hypothesises that, in the current context, the final set $\mathcal{A}$ has size $\left|{\mathcal{A}}\right|\geq(1-\rho)\left|{\mathcal{A}_{1}}\right|$, and likewise for $\mathcal{B}$, with $\rho=1/100$. Since the difference between $\mathcal{A}_{1}$ and $\mathcal{A}$ is just the $O(\epsilon_{D})n$ nodes deleted in Phase 2, this hypothesis is clearly satisfied, again without bothering to work out the small constant hidden in the $O(\cdot)$.
@@
Having controlled the size of $Q$ we are ready to consider the number of semirandom offers needed in Phase 2. In Phase 2a, that the first point offered is in $A_{0}$ occurs with probability just above $1/3$. The cardinality of $A$ is nearly that of $A_{0}$ and we have arranged that $\left|{Q}\right|\leq\frac{1}{100}\left|{A_{0}}\right|$, so $a\in A\setminus Q$ with probability at least
$4\displaystyle 1/3\cdot(1-O(\epsilon_{Q})-1/100)>1/$ 4 . (4.9)
@@
Independently (almost), the same holds for $a^{\prime}$, so (M2a-C1) holds with probability at least 1/161161/161 / 16.111We would have independence if the offered $a^{\prime}$ were independent of $a$, allowing the possibility that $a^{\prime}=a$ (in which case we would reject the offer). We can simulate this independence by tossing a coin and with probability $1/n$ replacing the offered $a^{\prime}$ with $a$. In the simulated model, (M2a-C1) and distinctness hold with probability at least $1/16-1/n$, or simply $1/16$ by absorbing the $1/n$ into slack in (4.9). Since the simulated substitution of $a^{\prime}$ for $a$ is never advantageous, (M2a-C1) holds with probability at least $1/16$ in the original model.
By 8, the chance that (M2a-C2) fails to hold is $O(1/n)$. Thus, each offer satisfies the conditions with probability at least $1/17$. Since at most $2\epsilon_{D}n$ actions are needed (see (4.4)), a.a.s., $18\cdot 2\epsilon_{D}n$ offers suffice.
@@
Similarly, in Phase 2b, an offer plus the added point $a^{\prime\prime}$ satisfies condition (M2b-C1) with probability more than $(1/4)^{3}$ (even if we chose $a^{\prime\prime}$ in [$n$] rather than $A$), and again (M2b-C2) fails with probability $O(1/n)$, so each offer results in action with probability at least $1/65$. By (4.8), the number of actions needed is at most $2\epsilon_{D}n$, so, a.a.s., $131\epsilon_{D}n$ offers suffice.
@@
Numerically, Phases 2a and 2b together require far under $0.001$n$ offers. More to the point, it is clear that, with control of $\left|{Q}\right|$, each offer results in action with probability $\Theta(1)$, and thus $O(n)$ offers suffice.
Recall that Phase 1 used $3Cn$ semirandom offers, with $C$ given by Lemma 6. Now that we have fixed $\epsilon_{D}=$0.000\hskip 1.30005pt001$$, the proof of Lemma 6 requires a $\lambda$ of 41, for $C$ a bit under $100$, so $300n$ semirandom offers suffice for Phase 1 (and the whole algorithm). We have made no effort whatsoever to optimise the constants.
@@ 4.4.2. Phase 2
For the stopping condition, since at the end of Phase 2a, $\left|{A}\right|\geq\left|{B}\right|$, and Phase 2b decreases the size of $A$ without changing $B$, it is clear that at some time we get $\left|{A}\right|\leq\left|{B}\right|$. At this time, we get equality if $\left|{A}\right|\equiv\left|{B}\right|\pmod{s}$, and this is so. Let $V=[n]$ be the set of all points, with $\left|{V}\right|\equiv 0\pmod{s}$. Let $\dot{V(D)}$ be the set of points in $D$. At this step, the points of $\dot{V(D)}$ are all in hyperedges in $\mathcal{M}$, so $\left|{V\setminus\dot{V(D)}}\right|\equiv 0\pmod{s}$. Thus, at this step,
@@ 4.4.3. Phase 3
As for matchings, it is enough to ensure that when constructing $D_{1}$, $\left|{X}\right|\leq\epsilon_{D}n$ for some sufficiently small constant $\epsilon_{D}$. The number of Phase 2 steps is $O(\epsilon_{D}n)$ (the $O(\cdot)$ expression hiding some constant independent of $\epsilon_{D}$), so the number of nodes deleted from $D$ is also $O(\epsilon_{D}n)$. By Lemma 9, $Q$ remains of size at most $k\>\eta_{c,k}(\delta)(O(\epsilon_{D}))n$ (plus $O(n^{2/3})$), and this can be made an arbitrarily small fraction of $n$ by choosing $\epsilon_{D}$ sufficiently small. If follows that each action is accepted w.p. $\Omega(1)$.

### v2

@@ 3. Outline, and lemmas on an auxiliary bipartite graph
Let $\rho=\frac{1-c}{c}\geq\max\left\{\frac{\left|{\mathcal{A}}\right|}{\left|{\mathcal{B}}\right|},\frac{\left|{\mathcal{B}}\right|}{\left|{\mathcal{A}}\right|}\right\}$. Consider a given set $S$ of size $\left|{S}\right|=\delta N$, fixed with knowledge of $\mathcal{A}$ and $\mathcal{B}$ but not the random edges. Each edge out of $B$ may be an inedge of $S\cap A$, so
$\operatorname{vol}(S\cap\mathcal{A})\sim\operatorname{Bin}\left(k\left|{B}\right|\;,\;{\left|{S\cap\mathcal{A}}\right|}/{\left|{\mathcal{A}}\right|}\,\right).$
The symmetric statement for $\operatorname{vol}(S\cap\mathcal{B})$ and the two volumes are independent, so $\operatorname{vol}(S)$ is a sum of independent Bernoulli random variables with expectation
$\mathbb{E}(\operatorname{vol}(S))$ $=k\left|{S\cap\mathcal{A}}\right|\,\tfrac{\left|{\mathcal{B}}\right|}{\left|{\mathcal{A}}\right|}+k\left|{S\cap\mathcal{B}}\right|\,\tfrac{\left|{\mathcal{A}}\right|}{\left|{\mathcal{B}}\right|}\;\leq k\;\left|{S}\right|\max\left\{\tfrac{\left|{\mathcal{A}}\right|}{\left|{\mathcal{B}}\right|},\tfrac{\left|{\mathcal{B}}\right|}{\left|{\mathcal{A}}\right|}\right\}\;\leq\;\rho k\delta N.$
For convenience we may artificially add more independent Bernoulli variables to obtain a random variable $Z$ stochastically dominating $\operatorname{vol}(S)$ and with $\mathbb{E}Z=\rho k\delta N$.
Let $\eta=\eta_{c,k}(\delta)$. Apply the Chernoff-type inequality $\Pr(Z\geq\mathbb{E}Z+t)\leq\exp\left(-\frac{t^{2}}{2(\mathbb{E}Z+t/3)}\right)$ [18, eq. (2.5) and Theorem 2.8] to $\Pr(Z\geq\eta N)$. This means taking $t=\eta N-\mathbb{E}Z=(r-1)\rho k\delta N$, where $r=\eta/(\rho k\delta)\geq 2$ by definition of $\eta_{c,k}(\delta)$. Then,
$\Pr(Z\geq\eta N)$ $\leq\exp\left(-\frac{(r-1)^{2}(\mathbb{E}Z)^{2}}{2(1+\frac{r-1}{3})\mathbb{E}Z}\right)\;\leq\;\exp\left(-\frac{(r/2)^{2}}{2(\frac{2r}{3})}\mathbb{E}Z\right)\;=\;\exp\Big{(}-\tfrac{3}{16}r\cdot\rho\delta kN\Big{)}.$
@@
$\binom{N}{\delta N}\Pr(Z\geq\eta N)$ $\leq\exp\Big{(}\delta N\cdot\Big{(}\ln(e/\delta)-\frac{3}{16}r\rho k\Big{)}\Big{)}$
@@
Given $k\geq 10$ and $\rho<1-2/(1+3/e)$. (Having $\rho\leq 0.049$ suffices.) Let $\mathcal{D}$ be a uniformly random $k$-out bipartite digraph on parts $\mathcal{A}$ and $\mathcal{B}$. Consider the subgraph $\mathcal{D}^{\prime}$ induced by a pair of subsets $\mathcal{A}^{\prime}\subseteq\mathcal{A}$, $\mathcal{B}^{\prime}\subseteq\mathcal{B}$, where
@@
Then, a.a.s. in $\left|{\mathcal{A}}\right|$ (with probability tending to 1 as $\left|{\mathcal{A}}\right|\rightarrow\infty$) every such induced subgraph $\mathcal{D}^{\prime}$ contains a perfect matching.
Note that here “a.a.s. in $\left|{\mathcal{A}}\right|$” is equivalent to a.a.s. in $\left|{\mathcal{B}}\right|$, $\left|{\mathcal{A}^{\prime}}\right|$, or $\left|{\mathcal{B}^{\prime}}\right|$. Also, the hypothesis $k\geq 10$ can be weakened to $k\geq 7$, at the expense of a slightly more complicated proof, and likely further.
@@
the last inequality follows from the hypothesis on $\rho$ and is used in the next calculation. The same holds for $\mathcal{B}$. Also, $\left|{\mathcal{A}}\right|$ and $\left|{\mathcal{B}}\right|$ are both $\Theta(n)$, and henceforth we work asymptotically in $n$.
@@ 4. Matching Strategy and Analysis
As a warmup exercise, consider the 1-offer model, and take the set sizes to be $\left|{A_{0}}\right|=\left|{B_{0}}\right|=n/3$. Construct a random 2-out bipartite graph as follows. Focussing on the first point $a_{1}\in A_{0}$, on offer of any $b\in\mathbf{b}\in B_{0}$, choose as third point the partner $b^{\prime}$ of $b$ in $\mathbf{b}$, construct the hyperedge $\left\{a_{1}\right\}\cup\mathbf{b}$, and correspondingly add the edge $\left\{a_{1},\mathbf{b}\right\}$ to the auxiliary graph. Repeat this to build a second random auxiliary graph edge on $a_{1}$, and the corresponding hyperedge; then repeat to build two edges on $a_{2}$, on $a_{3}$, and so forth. Similarly, focussing on the first point $\mathbf{b}_{1}\in B_{0}$, on offer of any $a\in A_{0}$, build the hyperedge $\left\{a,\mathbf{b}_{1}\right\}$ and the corresponding auxiliary graph edge $(a,\mathbf{b}_{1})$, repeating to make two such random edges (and corresponding hyperedges) on each node in $B_{0}$. This can be completed in linear time. (It will take about $2n$ trials to be offered $\frac{2}{3}n$ nodes in $A_{0}$ to make the requisite edges to $B_{0}$, and about $n$ trials to be offered $\frac{2}{3}n$ points in nodes of $B_{0}$ to make the edges to $A_{0}$, and these can even be done in parallel.) At this point the auxiliary graph is a random 2-out graph and thus a.a.s. has a perfect matching. By construction, the corresponding hyperedges form a hypergraph perfect matching, as desired.
In the 2-offer model we will be unable to make every node of $A_{0}$ and $B_{0}$ $k$-out, and will thus take the 3-phase approach outlined in Section 3.
@@ 4.2.1. Phase 2a
(Read the label as M for matching, 2a for the phase, and C for “condition”; in the next part, A is for “action”.)
@@ 4.2.3. Analysis and parameter choices
The number of inneighbours of these nodes is bounded by Lemma 9, which will be our focus for the next few paragraphs. The lemma’s $N$ is $\left|{A_{1}}\right|+\left|{B_{1}}\right|>\frac{2}{3}n$ by (4.2) and (4.3), so $10\epsilon_{D}n$ represents at most a $15\epsilon_{D}$ fraction of $N$. Thus we take the lemma’s $\delta$ — call it $\delta_{D}$ — equal to $15\epsilon_{D}$.
The two parts $A_{1}$ and $B_{1}$ are of nearly equal size, so we may take Lemma 9’s $c$ as $\frac{1}{3}$. We will make many assertions of this sort, so let us be more precise this one time. In this case, from (4.1), (4.2), and (4.3) we have $(\frac{1}{3}+3\epsilon_{D})n\leq\left|{A_{1}}\right|\leq\left|{A_{0}}\right|=(\frac{1}{3}+4\epsilon_{D})n$ and $(\frac{1}{3}-3\epsilon_{D})n\leq\left|{B_{1}}\right|\leq\left|{B_{0}}\right|=(\frac{1}{3}-2\epsilon_{D})n$. We will choose $\epsilon_{D}$ small enough that it is obvious that each of $A_{1}$ and $B_{1}$ has size at least $1/3$rd that of their union. More to the point, forgetting the precise multiples of $\epsilon_{D}$ here, we will have expressions like $\left|{A_{1}}\right|=(\frac{1}{3}+O(\epsilon_{D}))n$, and it will be clear that $\epsilon_{D}$ can be chosen sufficiently small to satisfy the various assertions.
Continuing, by Lemma 9, the number of inneighbours is at most $\eta_{1/3,k}(\delta_{D})N$. Each inneighbour has $k$ outneighbours, so the number of outneighbours is at most $k\>\eta_{1/3,k}(\delta_{D})N$. Since $\left|{Q_{1}}\right|=o(n)=o(N)$, we have that at all times $\left|{Q}\right|\leq k\>\eta_{1/3,k}(\delta)N+o(N)$. We will contrive that $k\>\eta_{1/3,k}(\delta_{D})\leq\tfrac{1}{10}$, so that $\left|{Q}\right|\leq\tfrac{N}{9}$ and $\left|{Q}\right|/\left|{A_{1}}\right|\leq(\tfrac{N}{9})/(\tfrac{N}{3})\leq\tfrac{1}{3}$ and symmetrically $\left|{Q}\right|/\left|{B_{1}}\right|\leq\tfrac{1}{3}$.
@@
This dictates our parameter choices. We take $k=10$ throughout, for purposes of Lemma 10. Then, in Lemma 9, choosing $\delta=\delta_{D}=$0.000\hskip 1.30005pt15$$ suffices, giving $k\>\eta_{1/3,k}(\delta_{D})\leq\tfrac{1}{10}$ as desired. (The value of $\eta$, defined in Lemma 9, is determined by the first term in the max; the second term, $40\>\delta$, is smaller.) Since we set $\delta_{D}=15\epsilon_{D}$, taking $\epsilon_{D}=$0.000\hskip 1.30005pt01$$ suffices.
[…]
Looking ahead, we will invoke Lemma 10 to prove existence of a perfect matching in $D$. The lemma’s $\mathcal{A}$ and $\mathcal{B}$ are the parts $A_{1}$ and $B_{1}$ at the start of Phase 2, its $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$ are the parts $A$ and $B$ at the end of Phase 2, and its $\rho$ is (an upper bound on) the fraction of nodes deleted, so we take $\rho=\delta$. With the choices above, then, $\rho\leq 0.049$, satisfying the hypothesis.
Having controlled the size of $Q$, we are ready to consider the number of semirandom offers needed in Phase 2.
@@
In Phase 2a, that the first point offered is in $A_{0}$ occurs with probability just above $1/3$. The cardinality of $A$ is nearly that of $A_{0}$ and we have arranged that $\left|{Q}\right|\leq\frac{1}{3}\left|{A_{1}}\right|$, so
$\Pr(a\in A\setminus Q)$ $\geq\tfrac{1}{3}\cdot(1-O(\epsilon_{D})-\tfrac{1}{3})\;>\;\tfrac{1}{5}.$ (4.9)
@@
Independently (almost), the same holds for $a^{\prime}$, so (M2a-C1) holds with probability at least 1/251251/251 / 25.222We would have independence if the offered $a^{\prime}$ were independent of $a$, allowing the possibility that $a^{\prime}=a$ (in which case we would reject the offer). We can simulate this independence by tossing a coin and with probability $1/n$ replacing the offered $a^{\prime}$ with $a$. In this simulated model, (M2a-C1) and distinctness hold with probability at least $\Pr(a\in A\setminus Q)^{2}-1/n$, here simply $1/25$ by absorbing the $1/n$ into slack in (4.9). Since the simulated substitution of $a^{\prime}$ for $a$ is never advantageous, (M2a-C1) holds with probability at least $1/25$ in the original model. By 8, the chance that (M2a-C2) fails to hold is $O(1/n)$. Thus, each offer satisfies the conditions with probability at least $1/26$. Since at most $2\epsilon_{D}n$ actions are needed (see (4.4)), a.a.s., $27\cdot 2\epsilon_{D}n$ offers suffice.
@@
Similarly, in Phase 2b, an offer plus the added point $a^{\prime\prime}$ satisfies condition (M2b-C1) with probability more than $(1/5)^{3}$ (even if we chose $a^{\prime\prime}$ in [$n$] rather than $A$), and again (M2b-C2) fails with probability $O(1/n)$, so each offer results in action with probability at least $1/126$. By (4.8), the number of actions needed is at most $2\epsilon_{D}n$, so, a.a.s., $127\cdot 2\epsilon_{D}n$ offers suffice.
@@
It is clear that, with control of $\left|{Q}\right|$, each offer in Phases 2a and 2b results in action with probability $\Theta(1)$, and thus $O(n)$ offers suffice. Numerically, Phases 2a and 2b together require at most $(27+127)\cdot 2\epsilon_{D}<$0.01$n$ offers, so $241n$ offers suffice for Phases 1 and 2 (thus for the whole algorithm). We have made no effort to optimise the constants.
@@ 4.4.2. Phase 2
For the stopping condition, since at the end of Phase 2a, $\left|{A}\right|\geq\left|{B}\right|$, and Phase 2b decreases the size of $A$ without changing $B$, it is clear that at some time we get $\left|{A}\right|\leq\left|{B}\right|$. At this time, we get equality if $\left|{A}\right|\equiv\left|{B}\right|\pmod{s}$, and this is so. Let $V=[n]$ be the set of all points, with $\left|{V}\right|\equiv 0\pmod{s}$. Let $\dot{V}(D)$ be the set of points in $D$, i.e., the points in $A$ and the points belonging to bases in $B$. At this step, the points of $V\setminus\dot{V}(D)$ are all in hyperedges in $\mathcal{M}$, so $\left|{V\setminus\dot{V}(D)}\right|\equiv 0\pmod{s}$. Thus, at this step,
@@ 4.4.3. Phase 3
As for $s=3$, it is enough to ensure that when constructing $D_{1}$, $\left|{X}\right|\leq\epsilon_{D}n$ for some sufficiently small constant $\epsilon_{D}$. The number of Phase 2 steps is $O(\epsilon_{D}n)$ (the $O(\cdot)$ expression hiding some constant independent of $\epsilon_{D}$), so the number of nodes deleted from $D$ is also $O(\epsilon_{D}n)$. By Lemma 9, $Q$ remains of size at most $k\>\eta_{c,k}(\delta)(O(\epsilon_{D}))n$ (plus $O(n^{2/3})$), and this can be made an arbitrarily small fraction of $n$ by choosing $\epsilon_{D}$ sufficiently small. If follows that each action is accepted w.p. $\Omega(1)$.